# AHA! Animating Human Avatars in Diverse Scenes with Gaussian Splatting

### v1

@@ Abstract
We present a novel framework for animating humans in 3D scenes using 3D Gaussian Splatting (3DGS), a neural scene representation that has recently achieved state-of-the-art photorealistic results for novel-view synthesis but remains under-explored for human-scene animation and interaction. Unlike existing animation pipelines that use meshes or point clouds as the underlying 3D representation, our approach introduces the use of 3DGS as the 3D representation to the problem of animating humans in scenes. By representing humans and scenes as Gaussians, our approach allows for geometry-consistent free-viewpoint rendering of humans interacting with 3D scenes. Our key insight is that the rendering can be decoupled from the motion synthesis and each sub-problem can be addressed independently, without the need for paired human-scene data. Central to our method is a Gaussian-aligned motion module that synthesizes motion without explicit scene geometry, using opacity-based cues and projected Gaussian structures to guide human placement and pose alignment. To ensure natural interactions, we further propose a human–scene Gaussian refinement optimization that enforces realistic contact and navigation. We evaluate our approach on scenes from Scannet++ and the SuperSplat library, and on avatars reconstructed from sparse and dense multi-view human capture. Finally, we demonstrate that our framework allows for novel applications such as geometry-consistent free-viewpoint rendering of edited monocular RGB videos with new animated humans, showcasing the unique advantage of 3DGS for monocular video-based human animation.
@@ 1 Introduction
Human animation in 3D scenes is essential for applications ranging from video gaming and computer-generated imagery (CGI) to robotics. Recent research has made significant progress on generating humans in 3D scenes Hassan et al. (2019); Jiang et al. (2024a); Hassan et al. (2021b); Zhao et al. (2023); Hwang et al. (2025), where humans are typically represented as either 3D skeletons or meshes, while background scenes are represented as meshes or point clouds. These representations are compact and versatile, capable of modeling a wide variety of surfaces. However, a fundamental limitation persists: there is almost always a domain gap between rendered results and real images due to limitations in lighting, materials, and geometric fidelity.
@@
In parallel, neural scene representations have emerged, beginning with NeRF Mildenhall et al. (2020) and recently 3D Gaussian Splatting Kerbl et al. (2023), enabling photorealistic rendering of objects, humans, and full 3D scenes from novel viewpoints or in novel poses. Yet, despite their success in rendering quality, neural representations have seen little to no adoption in human-scene animation pipelines, which continue to rely on mesh and point cloud–based frameworks.
@@
Gaussian Splatting as a 3D representation for human-scene animation, in theory offers natural advantages over existing mesh based representations: First, 3DGS enables photorealistic rendering of human-scene interactions with superior lighting and material fidelity. Second, 3DGS allows for reconstructing scenes with only a monocular video Ling et al. (2024) captured from a mobile phone, thus allowing for applications such as geometry consistent free viewpoint rendering of videos with new humans, personalized content creation and gaming in scenes from monocular mobile-captured videos. Such applications are difficult with a mesh based representation as estimating meshes, or pointclouds from only monocular scene videos remains challenging Wang et al. (2025). This motivates the central question addressed in this paper: Can neural scene representations—specifically Gaussian Splatting—be effectively used as a 3D representation for human animation in 3D scenes? (Fig. LABEL:fig:teaser).
@@
Several obstacles prevent a direct extension of 3DGS to human animation in 3D scenes. First, most existing work on human–scene interaction synthesis Hassan et al. (2021b); Jiang et al. (2024a); Hwang et al. (2025); Zhao et al. (2023) assumes paired human motion data with scene geometry. Such datasets are difficult to collect at scale, and conversion from mesh-based annotations into Gaussians is non-trivial. Second, human-scene animation requires motion synthesis that respects both the structure of the scene and the natural dynamics of the human pose, which for a non-mesh representation remains non-trivial. Furthermore, unlike meshes, 3DGS does not provide explicit topology or clean geometry, complicating tasks like surface-based contact modeling.
@@
To address these challenges, we offer a novel perspective for human–scene animation, grounded in two key insights: First, the rendering of humans and scenes in 3DGS can be decoupled from motion synthesis. That is, we can reconstruct humans and scenes independently, animate humans in a canonical space, and then place them back into reconstructed 3DGS scenes. This philosophy is common in classical graphics pipelines for meshes, where canonical models are animated via skinning or rigging, and has recently been adopted for animatable 3DGS avatars as well. However, prior work has primarily studied such Gaussian avatars in isolation. In contrast, our contribution is to extend this paradigm to human–scene animation, where avatars must not only be animated but also consistently placed and rendered inside reconstructed 3DGS scenes. Second, motion synthesis can itself be decoupled from explicit geometry: even though 3DGS does not provide watertight surfaces, we show that its opacity fields and projected Gaussian structures offer sufficient cues to guide human placement.
@@
Building on these insights, our framework proceeds in two stages. First, we reconstruct humans as animatable Gaussians from either sparse or dense multiview capture using an off-the-shelf module. These Gaussians are then posed using a Gaussian-aligned motion module (Sec. 3.2), which computes scene-aligned motion parameters by relying on opacity-based culling and orthographic projection of scene Gaussians for path finding. This allows us to place reconstructed humans into new 3DGS scenes in a geometrically consistent manner. To further ensure realistic interactions, we introduce a human–scene Gaussian refinement optimization (Sec. 3.3) that adjusts the placement and motion of humans for natural contact and navigation within the scene.
@@
To showcase the wide-applicability of our method on diverse datasets, we present results on scenes from the Scannet++ dataset Yeshwanth et al. (2023) and from scenes downloaded from the publicly available SuperSplat 3DGS library SuperSplat . To demonstrate the efficacy of our method on different Avatar reconstruction datasets, we also showcase results on Avatars from the BEHAVE Bhatnagar et al. (2022) and Avatarrex datasets Zheng et al. (2023). The results from BEHAVE demonstrate that our method works on avatars reconstructed from sparse camera setups. We finally demonstrate the utility of our presented framework for geometry consistent free viewpoint rendering of monocular videos with new animated humans on several monocular videos from the DL3DV dataset Ling et al. (2024), showcasing the unique advantage of 3DGS for casual video-based human animation.
@@ 2 Related Work
Neural Rendering Following the publication of NeRF Mildenhall et al. (2020), there has been significant research on Neural Rendering  Xie et al. (2022b). Nerf is limited by its computational complexity and despite several follow-up improvements Müller et al. (2022); Barron et al. (2022; 2023); Tancik et al. (2023), the high computational cost of NeRF rermain. 3DGS introduced in Kerbl et al. (2023) addresses this limitation by representing scenes with an explicit set of primitives shaped as 3D Gaussians, extending previous work  Lassner & Zollhöfer (2021). 3DGS rasterizes Gaussian primitives into images using a splatting algorithm Westover (1992). 3DGS originally designed for static scenes has been extended to dynamic scenes  Shaw et al. (2023); Luiten et al. (2024); Wu et al. (2024); Lee et al. (2024); Li et al. (2023a), slam-based reconstruction, Keetha et al. (2024), mesh reconstruction Huang et al. (2024); Guédon & Lepetit (2024) and NVS from sparse cameras Mihajlovic et al. (2024).
@@
Human Reconstruction and Neural Rendering Mesh-based templates Pavlakos et al. (2019); Loper et al. (2015) have been used to recover 3D human shape and pose from images and video Bogo et al. (2016); Kanazawa et al. (2018). However, this does not allow for photoreal renderings. In Alldieck et al. ; Alldieck et al. (2019) recover a re-posable human avatar from monocular RGB. However their use of a mesh template also does not allow for photorealistic renderings. Implicit functions Mescheder et al. (2019); Park et al. (2019) have also been utilized to reconstruct detailed 3D clothed humans Chen et al. (2021); Alldieck et al. (2021); Saito et al. (2020); He et al. (2021); Huang et al. (2020); Deng et al. (2020). However, they are also unable to generate photorealistic renderings and are often not reposable. Several works Peng et al. (2021); Guo et al. (2023); Weng et al. (2022); Jiang et al. (2022); Habermann et al. (2023); Zhu et al. (2024); Li et al. (2022); Liu et al. (2021); Xu et al. (2021) build a controllable NeRF that produces photorealistic images of humans from input videos. Unlike us, they do not model human-scene interactions. With the advent of 3DGS, several recent papers use a 3DGS formulation Kocabas et al. (2023); Qian et al. (2024); Moreau et al. (2024); Abdal et al. (2024); Zielonka et al. (2023); Moon et al. (2024); Li et al. (2024b); Pang et al. (2024); Lei et al. (2023); Hu et al. (2024); Li et al. (2024a); Zheng et al. (2024); Jiang et al. (2024b); Dhamo et al. (2024); Qian et al. (2023); Xu et al. (2024); Junkawitsch et al. (2025) to build controllable human or face avatars. Unlike our method, they do not model human-scene interactions. Prior works have also extended the 3DGS formulation to model humans along with their environment, Xue et al. (2024); Zhan et al. (2024); Mir et al. (2025), but unlike us, they either do not focus on human animation in 3D scenes.
@@
Humans and Scenes Human-scene interaction is a recurrent topic of study in computer vision and graphics. Early works Fouhey et al. (2014); Wang et al. (2017); Gupta et al. (2011) model affordances and human-object interactions using monocular RGB. The collection of several recent human-scene interaction datasets Hassan et al. (2021a); Guzov et al. (2021); Hassan et al. (2019); Savva et al. (2016); Taheri et al. (2020); Bhatnagar et al. (2022); Jiang et al. (2024a); Cheng et al. (2023); Zhang et al. (2022) has allowed the computer vision community to make significant progress in joint 3D reconstruction of human-object interactions Xie et al. (2022a; 2023; 2024a); Zhang et al. (2020). These datasets have also led to the development of methods that synthesize object conditioned controllable human motion Zhang et al. (2022); Starke et al. (2019b); Hassan et al. (2021c); Diller & Dai (2024). All these methods represent humans and scenes as 3D meshes and inherit the limitations of mesh-based representations including their inability to generate photorealistic images, while our method allows for photorealistic renderings of humans and scenes.
@@
Human Animation Human animation is another extensively studied problem in vision and graphics. Motion matching (Reitsma & Pollard, 2007), learned motion matching (Clavet, 2016; Holden et al., 2020) and motion graphs  (Lee et al., 2002; Fang & Pollard, 2003; Kovar et al., 2008; Safonova et al., 2004; Safonova & Hodgins, 2007) are common methods employed in the video-gaming industry for generating kinematic motion sequences. Deep learning variants Holden et al. (2017); Starke et al. (2019a; 2021; 2020) have also gained popularity. Diffusion Models Tevet et al. (2023) have emerged as a powerful paradigm for human motion synthesis. Several follow-up works extend the original Motion Diffusion model with physicsYuan et al. (2023), blended-positional encoding Barquero et al. (2024), and for fine-grained controllable motion synthesis Karunratanakul et al. (2023); Pinyoanuntapong et al. (2024); Xie et al. (2024b). Reinforcement learning Zhang & Tang (2022); Zhao et al. (2023) is another oft-used paradigm used for motion synthesis. Diffusion models have also been used as latent-motion models Zhao et al. (2025) but unlike us, they only focus on clean, noise-free mesh-based scene representations and ignore the rendering aspects of human-scene interaction leading to limited rendering quality and photorealism.
@@
Figure 1: Method: Top Left: Using multiview images of the human we reconstruct controllable human Gaussians. Bottom Left: Using scene video we reconstruct static scene Gaussians. Top Middle: We synthesize human motion that confirms with the 3D scene Gaussians using Gaussian-aligned motion module. Top Right: The human poses are used to animate human Gaussians. We further refine these Gaussians for correct placements and contact. Bottom Right: These composited human and scene Gausssians can be rendered from any viewpoint to generate photoreal images of human-scene interaction.
@@ 3 Method
We present a method that enables virtual humans represented using Gaussian splats to navigate and interact in complex environments reconstructed as 3D Gaussian scenes. Our framework consists of three key components: (1) Gaussian Reconstruction: We reconstruct both scenes and humans as 3D Gaussians from RGB images. For scenes, we use standard 3DGS reconstruction, while for humans, we learn human Gaussian representations that can be animated with different SMPL poses (Sec. 3.1). (2) Gaussian-Aligned Motion Synthesis: Central to our approach is a novel Gaussian-aligned motion module (Sec. 3.2), which uses the reconstructed scenes (Sec. 3.1) and a latent-variable based motion synthesis framework (using RL and latent space optimization) to synthesize motion parameters aligned with the 3DGS scenes (3) Differentiable Contact Refinement in 3DGS: We use the synthesized human motion data to animate human Gaussians and apply a novel refinement algorithm for correct human-scene interaction (Sec. 3.3). The refinement module detects contact frames from motion data and optimizes translation vectors to enforce proper contact between human and scene Gaussians while maintaining temporal smoothness and avoiding penetrations. These composed Gaussians can be rendered from any camera viewpoint to produce videos of humans interacting with diverse scenes. Figure 1 provides an overview.

### v2

@@ Abstract
We present a novel framework for animating humans in 3D scenes using 3D Gaussian Splatting (3DGS), a neural scene representation that has recently achieved state-of-the-art photorealistic results for novel-view synthesis but remains under-explored for human-scene animation and interaction. Unlike existing animation pipelines that use meshes or point clouds as the underlying 3D representation, our approach introduces the use of 3DGS as the 3D representation to the problem of animating humans in scenes. By representing humans and scenes as Gaussians, our approach allows for geometry-consistent free-viewpoint rendering of humans interacting with 3D scenes. Our key insight is that the rendering can be decoupled from the motion synthesis and each sub-problem can be addressed independently, without the need for paired human-scene data. Central to our method is a Gaussian-aligned motion module that synthesizes motion without explicit scene geometry, using opacity-based cues and projected Gaussian structures to guide human placement and pose alignment. To ensure natural interactions, we further propose a human–scene Gaussian refinement optimization that enforces realistic contact and navigation. We evaluate our approach on scenes from Scannet++ and the SuperSplat library, and on avatars reconstructed from sparse and dense multi-view human capture. Finally, we demonstrate that our framework allows for novel applications such as geometry-consistent free-viewpoint rendering of edited monocular RGB videos with new animated humans, showcasing the unique advantage of 3DGS for monocular video-based human animation. To gauge the full quality of our results we urge the reader to watch the supplementary available at https://miraymen.github.io/aha/
@@ 1 Introduction
Human animation in 3D scenes is essential for applications ranging from video gaming and computer-generated imagery (CGI) to robotics. Recent research has made significant progress on generating humans in 3D scenes (hassan2019prox; jiang2024scaling; hassan_samp_2021; zhao2023dimos; hwang2025scenemimotioninbetweeningmodeling), where humans are typically represented as either 3D skeletons or meshes, while background scenes are represented as meshes or point clouds. These representations are compact and versatile, capable of modeling a wide variety of surfaces. However, a fundamental limitation persists: there is almost always a domain gap between rendered results and real images due to limitations in lighting, materials, and geometric fidelity.
@@
In parallel, neural scene representations have emerged, beginning with NeRF (mildenhall2020nerf) and recently 3D Gaussian Splatting (kerbl3Dgaussians), enabling photorealistic rendering of objects, humans, and full 3D scenes from novel viewpoints or in novel poses. Yet, despite their success in rendering quality, neural representations have seen little to no adoption in human-scene animation pipelines, which continue to rely on mesh and point cloud–based frameworks.
@@
Gaussian Splatting as a 3D representation for human-scene animation, in theory offers natural advantages over existing mesh based representations: First, 3DGS enables photorealistic rendering of human-scene interactions with superior lighting and material fidelity. Second, 3DGS allows for reconstructing scenes with only a monocular video (ling2024dl3dv) captured from a mobile phone, thus allowing for applications such as geometry consistent free viewpoint rendering of videos with new humans, personalized content creation and gaming in scenes from mobile-captured videos. Such applications are difficult with a mesh based representation as estimating meshes, or pointclouds from only monocular scene videos remains challenging (wang2025vggt). This motivates the central question addressed in this paper: Can neural scene representations—specifically Gaussian Splatting—be effectively used as a 3D representation for human animation in 3D scenes? (Fig. LABEL:fig:teaser).
@@
Several obstacles prevent a direct extension of 3DGS to human animation in 3D scenes. First, most existing work on human–scene interaction synthesis (hassan_samp_2021; jiang2024scaling; hwang2025scenemimotioninbetweeningmodeling; zhao2023dimos) assumes paired human motion data with scene geometry. Such datasets are difficult to collect at scale, and conversion from mesh-based annotations into Gaussians is non-trivial. Second, human-scene animation requires motion synthesis that respects both the structure of the scene and the natural dynamics of the human pose, which for a non-mesh representation remains non-trivial. Furthermore, unlike meshes, 3DGS does not provide explicit topology or clean geometry, complicating tasks like surface-based contact modeling.
@@
To address these challenges, we offer a novel perspective for human–scene animation, grounded in two key insights: First, rendering of humans and scenes in 3DGS can be decoupled from motion synthesis. That is, we can reconstruct humans and scenes independently, animate humans in a canonical space, and then place them back into reconstructed 3DGS scenes. This is common in classical graphics pipelines for meshes, where canonical models are animated via skinning or rigging, and has recently been adopted for animatable 3DGS avatars as well. However, prior work has primarily studied such Gaussian avatars in isolation. In contrast, our contribution is to extend this paradigm to human–scene animation, where avatars must not only be animated but also consistently placed and rendered inside reconstructed 3DGS scenes. Second, motion synthesis can itself be decoupled from explicit geometry: even though 3DGS does not provide watertight surfaces, we show that its opacity fields and projected Gaussian structures offer sufficient cues to guide human placement.
@@
Our framework proceeds in two stages. First, we reconstruct humans as animatable Gaussians from either multiview capture using an off-the-shelf module. These Gaussians are then posed using a Gaussian-aligned motion module (Sec. 3.2), which computes scene-aligned motion parameters by relying on opacity-based culling and orthographic projection of scene Gaussians for path finding. Our core contribution is the migration and adaptation of traditional scene-mesh human interaction techniques (including RL-based locomotion and motion transitions) to 3DGS. To further ensure realistic interactions, we introduce a human–scene Gaussian refinement optimization (Sec. 3.3) that adjusts the placement and motion of humans for natural contact and navigation within the scene.
@@
To showcase the applicability of our method on diverse datasets, we present results on scenes from the Scannet++ dataset (yeshwanthliu2023scannetpp) and from scenes downloaded from the publicly available SuperSplat 3DGS library (supersplat). To demonstrate the efficacy of our method on different Avatar reconstruction datasets, we also showcase results on Avatars from the BEHAVE (bhatnagar22behave) and Avatarrex datasets (zheng2023avatarrex). The results from BEHAVE demonstrate that our method works on avatars reconstructed from sparse camera setups. We finally demonstrate the utility of our presented framework for geometry consistent free viewpoint rendering of monocular videos with new animated humans on several monocular videos from the DL3DV dataset (ling2024dl3dv), showcasing the unique advantage of 3DGS for casual video-based human animation.
@@ 2 Related Work
Neural Rendering Following the publication of NeRF (mildenhall2020nerf), there has been significant research on Neural Rendering  (nerf_review). Nerf is limited by its computational complexity and despite several follow-up improvements (mueller2022instant; barron2022mip; barron2023zipnerf; nerfstudio), the high computational cost of NeRF rermain. 3DGS introduced in (kerbl3Dgaussians) addresses this limitation by representing scenes with an explicit set of primitives shaped as 3D Gaussians, extending previous work  (lassner2021pulsar). 3DGS rasterizes Gaussian primitives into images using a splatting algorithm (westover1991phdsplatting). 3DGS originally designed for static scenes has been extended to dynamic scenes  (shaw2023swags; luiten2023dynamic; Wu2024CVPR; lee2024ex4dgs; li2023spacetime), slam-based reconstruction, (keetha2024splatam), mesh reconstruction (Huang2DGS2024; guedon2023sugar) and NVS from sparse cameras (mihajlovic2024SplatFields).
@@
Human Reconstruction and Neural Rendering Mesh-based templates (SMPL-X:2019; smpl2015loper) have been used to recover 3D human shape and pose from images and video (Bogo2016keepitsmpl; kanazawaHMR18). However, this does not allow for photoreal renderings. In (alldieck2018video; alldieck19cvpr) recover a re-posable human avatar from monocular RGB. However their use of a mesh template also does not allow for photorealistic renderings. Implicit functions (mescheder2019occupancy; park2019deepsdf) have also been utilized to reconstruct detailed 3D clothed humans (chen2021snarf; alldieck2021imghum; saito2020pifuhd; he2021arch++; huang2020arch; deng2020nasa). However, they are also unable to generate photorealistic renderings and are often not reposable. Several works (peng2021animatable; guo2023vid2avatar; weng_humannerf_2022_cvpr; jian2022neuman; haberman2023hdhumans; heminggaberman2024trihuman; li2022tava; liu2021neural; xu2021hnerf) build a controllable NeRF that produces photorealistic images of humans from input videos. Unlike us, they do not model human-scene interactions. With the advent of 3DGS, several recent papers use a 3DGS formulation (kocabas2023hugs; qian20233dgsavatar; moreau2024human; abdal2023gaussian; zielonka2023drivable; moon2024exavatar; li2024animatablegaussians; pang2024ash; lei2023gart; hu2024gaussianavatar; li2024animatable2; zheng2024gpsgaussian; jiang2024robust; dhamo2023headgas; qian20233dgs; xu2024gaussian; junkawitsch2025eva) to build controllable human or face avatars. Unlike our method, they do not model human-scene interactions. Prior works have also extended the 3DGS formulation to model humans along with their environment, (xue2024hsr; tomie; mir2025gaspacho), but unlike us, they either do not focus on human animation in 3D scenes.
@@
Humans and Scenes Human-scene interaction is a recurrent topic of study in computer vision and graphics. Early works (fouhey2014people; wang2017binge; gupta20113d) model affordances and human-object interactions using monocular RGB. The collection of several recent human-scene interaction datasets (Hassan2021-gb; mir20hps; hassan2019prox; savva2016pigraphs; taheri2020grab; bhatnagar22behave; jiang2024scaling; cheng2023dnarendering; zhang2022couch) has allowed the computer vision community to make significant progress in joint 3D reconstruction of human-object interactions (xie2022chore; xie2023vistracker; xie2024template; zhang2020phosa). These datasets have also led to the development of methods that synthesize object conditioned controllable human motion (zhang2022couch; starke2019neuralstate; hassan21cvpr; diller2023cghoi). All these methods represent humans and scenes as 3D meshes and inherit the limitations of mesh-based representations including their inability to generate photorealistic images, while our method allows for photorealistic renderings of humans and scenes.
@@
Human Animation Human animation is another extensively studied problem in vision and graphics. Motion matching (reitsma2007evaluating), learned motion matching (clavet16motionmatching; holden2020learned) and motion graphs  (lee2002interactive; fang2003efficient; kovar2008motion; safonova2004synthesizing; Safonova:2007:InterpolatedGraphs) are common methods employed in the video-gaming industry for generating kinematic motion sequences. Deep learning variants (holden2017phase; starke19neural; starke21martialarts; starke20local) have also gained popularity. Diffusion Models (tevet2023human) have emerged as a powerful paradigm for human motion synthesis. Several follow-up works extend the original Motion Diffusion model with physics(yuan2023physdiff), blended-positional encoding (barquero2024seamless), and for fine-grained controllable motion synthesis (karunratanakul2023dno; pinyoanuntapong2024controlmm; xie2024omnicontrol). Reinforcement learning (zhang2022wanderings; zhao2023dimos) is another oft-used paradigm used for motion synthesis. Diffusion models have also been used as latent-motion models (Zhao:DartControl:2025) but unlike us, they only focus on clean, noise-free mesh-based scene representations and ignore the rendering aspects of human-scene interaction leading to limited rendering quality and photorealism.
@@
Figure 1: Using a latent motion model and 3DGS scene representation, we synthesize human motion that confirms with the 3D scene Gaussians using Gaussian-aligned motion module. We adapt RL based controllers and latent optimization for 3DGS scenes. We further refine these Gaussians for correct placements and contact. These composited human and scene Gausssians can be rendered from any viewpoint to generate photoeal images.
@@ 3 Method
We present a method that enables virtual humans represented using Gaussian splats to navigate and interact in complex environments reconstructed as 3D Gaussian scenes. Our framework consists of three key components: (1) Gaussian Reconstruction: We reconstruct both scenes and humans as 3D Gaussians from RGB images. For scenes, we use standard 3DGS reconstruction, while for humans, we learn human Gaussian representations that can be animated with different SMPL poses (Sec. 3.1). (2) Gaussian-Aligned Motion Synthesis: Central to our approach is a novel Gaussian-aligned motion module (Sec. 3.2), which uses the reconstructed scenes (Sec. 3.1) and a latent-variable based motion synthesis framework (using RL and latent space optimization adapted to 3DGS) to synthesize motion parameters aligned with the 3DGS scenes (3) Differentiable Contact Refinement in 3DGS: We use the synthesized human motion data to animate human Gaussians and apply a novel refinement algorithm for correct human-scene interaction (Sec. 3.3) (Fig. 2). The refinement module detects contact frames from motion data and optimizes translation vectors to enforce proper contact between human and scene Gaussians while maintaining temporal smoothness and avoiding penetrations. These composed Gaussians can be rendered from any camera viewpoint to produce videos of humans interacting with diverse scenes. Figure 1 provides an overview.